Adam is twice as old as Barry, Charlie is three years younger than Barry. The sum of their ages is 53, how old are they.
Answer by addingup(3677)

A = 2B
C = B-3
We will use these values for A and C in the next equation:
A+B+C = 53
2B+B-3+B = 53 add B on left and add 3 on both sides
4B = 56 divide both sides by 4
B = 14 And the problem says:
A = 2B = 2*14 = 28
C = B-3 = 14-3 = 11
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
14+28+11 = 53 Correct.
J